Find the equation if the line satisfying the given conditions, giving it in slope-intercept form if possible. Through (-4,-5), parallel to 6x + 5y= -34
Answer by stanbon(75887)

Find the equation if the line satisfying the given conditions, giving it in slope-intercept form if possible. Through (-4,-5), parallel to 6x + 5y= -34
---
Find the slope of the given line:
5y = -6x-34
y = (-6/5)x - (34/5)
---
slope = -6/5
----------------
Find equation of line passing thru (-4,-5) with slope = -6/5
---
Form: y = mx + b
-5 = (-6/5)(-4) + b
-5 = 24/5 + b
b = -49/5
-----
Equation: y = (-6/5)x - (49/5)
==================================
Cheers,
Stan H.
==========